OpenStudy (james1107):

Identify all the points that are on the following quadratic function: y = x2 - 3x (0,3) (0,0) (2,-2) (-1,4) (-3,15)

OpenStudy (james1107):

plz help will medal

OpenStudy (james1107):

can you explain this to me

OpenStudy (flower2468):

I am sorry, but I don't know anything about this, I can try to find someone who does

OpenStudy (james1107):

ok thanks

OpenStudy (james1107):

how bout @mathmale

OpenStudy (flower2468):

@ChrisCarmack @alivejeremy @-SpencerBraz- @fomf

OpenStudy (flower2468):

he would be a good idea to try

OpenStudy (james1107):

@Liz_Beauty45

OpenStudy (alivejeremy):

Still need help ?

OpenStudy (james1107):

yes plz

OpenStudy (alivejeremy):

The answer to this question is (0,0)

OpenStudy (james1107):

thank you sooooooooo much

OpenStudy (alivejeremy):

Tell me if it's wrong or right

OpenStudy (james1107):



OpenStudy (alivejeremy):

Your Welcome james

OpenStudy (james1107):

did you come up with more than one point

OpenStudy (alivejeremy):

Nope

OpenStudy (james1107):

oh well i think this had more than one answer go reread the question

OpenStudy (kayders1997):

Okay plug your x part into the fiction and see if it comes out to be the y part for the function

OpenStudy (james1107):

iv got the answer thanks everyone

OpenStudy (anonymous):

\[y = x2 - 3x \] \[ x=y^2-3\] \[x+3=y^2\]

OpenStudy (alivejeremy):

Yeah all i got was (0,)

OpenStudy (alivejeremy):

(0,0)*

OpenStudy (james1107):

lol

OpenStudy (anonymous):

wrong question

OpenStudy (kayders1997):

So I'll show you the first example you have the point (0,3) and x^2-3x so you x part of the coordinate is 0 so plug that into your equation (0)^2-3(0)=3? That is not true because 0 does not equal 3 thati S what you need to do for all these

OpenStudy (kayders1997):

There is more than one coordinate here though

OpenStudy (kayders1997):

@james1107 I found 3 correct

OpenStudy (james1107):

dont worry about it i alreadt turned it in

OpenStudy (kayders1997):

Okay